### v1

@@ 3.2 GPT’s coding ability on real and synthetic data
However, the performance on the MIMIC-IV test set is notably low, especially in precision. The improvement in the precision from leaf-only results to hierarchical is minimal. This implies that incorrect predictions were more likely to be out-of-family. Moreover, results on CoPHE are lower than on the set-based hierarchical evaluation indicating a tendency of the model to over-/under-predict within the scope of the family – an issue previously reported in local ICD coding models [32].
These results demonstrate that G can identify ICD-10 codes based on provided descriptions if presented within the prompts. Its performance when tasked with standard ICD coding without explicitly identified concepts or non-standard surface forms of the concepts significantly deteriorates.
@@
Real 9.46 33.511 14.755 10.585 44.873 17.129 10.3 44.33 16.715
Synthetic 59.064 40.715 48.203 66.461 41.319 50.958 67.198 41.545 51.346
Table 4: Results of GPT’s coding ability on real and self-generated data.
@@ 3.3 Acceptability of Generated Data in Clinical Practice
We calculated the inter-evaluator agreement for the 7 metrics using Fleiss’ kappa ($\kappa$) [33]. As $\kappa$ is designed for categorical variables and does not fully capture ordinal scores, we present it alongside the mean scores for each metric ($\mu$). The evaluators’ agreement was poor ($\kappa<0$) in examples from MIMIC-IV for Correctness and Informativeness of non-low-resource codes, and the Acceptability of the discharge summaries. For the other metrics, a $\kappa>0$ was reached but never exceeded 0.4 (lower than moderate agreement). All mean scores are higher than 4. Hence, while the clinicians disagreed on the exact scores, they rated real discharge summaries positively. The disagreement may be due to clinicians being UK-based with significant differences in reporting style within the UK and the US (where MIMIC-IV is from).
For GPT-generated summaries, slight agreement was seen in acceptability, and fair agreement in the correctness of low-resource labels. All other metrics had poor agreement. Both correctness metrics scored above 4, with low-resource correctness surpassing 4.5—an encouraging outcome for our primary goal of generating low-resource code data. Mean informativeness in the low-resource scenario and authenticity of scores were at least 3. Once again, performance on the low-resource codes exceeded non-low-resource codes. Other metrics had $\mu$ scores above 2. Informativeness and authenticity for non-low-resource codes had poor agreement ($\kappa<0$), while acceptability had some agreement with the lowest mean score of 2.225. The full table of $\kappa$ and $\mu$ scores for the evaluation can be found in Supplementary Material.
While G exhibits a good level of correctness as affirmed by the clinicians, several challenges persist in the generation of natural-looking clinical notes:
@@
G tends to generate clinical notes which list all concepts mentioned in the prompt. While an LLM following instructions is desirable behaviour, it is also important to exclude non-crucial details when tasked to generate a natural-looking clinical note. Real clinical notes often omit irrelevant normal findings and certain less-critical diagnoses within the context of an admission inferrable from supporting information such as medications and treatments for brevity. For instance, G unnecessarily noted a BMI within a normal range.
@@
G may also express diagnoses in technical and verbatim manners, not typical in clinical text. For instance, G mentioned “anaemia, which was unspecified.” in the generated clinical note as it was prompted with “D64.9: Anemia, unspecified”. G also occasionally introduced vague phrases such as “geriatric team provided supportive care, including behavioural interventions and medication management” without further detail. This unnatural style easily differentiates synthetic clinical notes from real ones.
3.3.3 GPT-3.5 misses necessary supporting information
G tends to produce clinical notes lacking crucial supporting information. For instance, G mentioned “Following a traumatic event” without specifying the event, which is unacceptable in the clinical setting. G also mentioned the introduction of pleural intervention without additional detail or rationale. This omission limits the overall informativeness of the patient’s medical context, potentially hindering the notes’ usability for a comprehensive view.
@@
The clinicians also observed that G sometimes introduces improbable but possible details. For instance, G overemphasised the significance of a patient’s anxiety disorder regarding an episode of syncope and a subsequent facial fracture, which the clinicians consider unlikely.
@@
G also failed to link diagnoses as interconnected events, crucial for coherent patient summaries. The clinicians described G-generated clinical notes as collections of unrelated facts. For example, G listed multiple complications of Type 1 diabetes mellitus (i.e., “E10.43”, “E10.621”, “E10.628”, “E10.21”, “E10.319”) separately, failing to acknowledge their connectedness. Within the same note G failed to recognise the connection between vascular complications and diabetes, presenting them as independent facts. The lack of coherence between diagnoses may impede the plausibility of the clinical note and undermine the overall acceptability and usefulness of synthetic notes.
@@
Furthermore, G failed at prioritising critical diagnoses, which undermined the authenticity of the portrayed scenario. For example, G presented multiple fractures alongside lesser issues, such as impacted ear wax, cataracts, and conjunctival haemorrhage. Hence, we concluded that G struggles to effectively convey the relative clinical significance of certain diagnoses.
@@ 4 DISCUSSION, CONCLUSION AND FUTURE STUDIES
In this work, we have investigated the capability of GPT-3.5’s potential in augmenting ICD-10 coding for local neural models in low-resource scenarios. While overall performance dipped with synthetic data augmentation, filtered codeset evaluation showed improvements, especially in advanced models like LAAT and Multi-Res CNN. Error analysis indicated augmented models made fewer out-of-family predictions, with some shift to within-family errors (closer to the correct answer). Augmentation showed promise in improving the prediction of generated codes and their siblings. Zero-shot labels did not consistently benefit from the augmentation, emphasising the need for real data in augmentation success. However, a zero-shot code learned from the synthetic data was predicted correctly.
In guided synthetic settings with ICD-10 descriptions, GPT-3.5 showed partial code identification ability displaying lesser over-/under-prediction tendencies than previously reported local models. It, however, struggled in the realistic scenario without in-prompt aid, performing below locally-trained models. Hence, it is unsuitable for deployment in a clinical setting.
@@
One potential solution to generating synthetic discharge summaries involves restructuring the prompt to order diagnoses chronologically, providing their corresponding timestamps. This could guide LLMs in creating synthetic notes mirroring chronological progression of a patient’s medical journey, enhancing coherence and prioritisation.
@@ 7 Author Contributions
MF proposed exploring data augmentation for ICD coding with GPT-3.5. and developed the main ideas of the publication with AG, HD, AB, and BA. MF, AG, HD, AB, and BA contributed to the overall experimental design. MF designed and performed the synthesis of new discharge summaries, and the evaluation of GPT’s ICD-10 coding performance on GPT-generated data. AG evaluated GPT’s ICD-10 coding performance on MIMIC-IV data. MF, AG, HD, and LD designed the evaluation of real versus generated discharge summaries by clinical staff. LD, SB, MH, RSP performed the clinical expert evaluation of the data. MF and AG analysed the results of the evaluation by clinicians. MF, AG, HD wrote the manuscript. All authors participated in editing the manuscript. BA and AB supervised the project and provided feedback and input on the experiments and analyses.

### v2

@@ 3.2 GPT’s coding ability on real and synthetic data
However, the performance on the MIMIC-IV test set is notably low, especially in precision. The improvement in the precision from leaf-only results to hierarchical is minimal. This implies that incorrect predictions were more likely to be out-of-family. Moreover, results on CoPHE are lower than on the set-based hierarchical evaluation indicating a tendency of the model to over-/under-predict within the scope of the family – an issue previously reported in local ICD coding models [32], and present in the reported hierarchical results for baseline LAAT.
These results demonstrate that G can identify ICD-10 codes in self-generated mentions based on provided descriptions if presented within the prompts. Its performance when tasked with standard ICD coding without explicitly identified concepts or non-standard surface forms of the concepts significantly deteriorates.
Table 3: Results of GPT-3.5’s coding ability on real and self-generated data.
@@
GPT-3.5 Real 9.46 33.51 14.76 10.59 44.87 17.13 10.30 44.33 16.72
GPT-3.5 Synthetic 59.06 40.72 48.20 66.46 41.32 50.96 67.20 41.55 51.35
Best Baseline (LAAT) Real 60.42 54.46 57.29 61.28 54.50 57.68 60.84 54.33 57.39
@@ 3.3 Acceptability of Generated Data in Clinical Practice
We calculated the inter-evaluator agreement for the 7 metrics using Fleiss’ kappa ($\kappa$) [33]. As $\kappa$ is designed for categorical variables and does not fully capture ordinal scores, also produced the mean scores for each metric ($\mu$). The results are presented in Table 4.
The evaluators’ agreement was poor ($\kappa<0$) in examples from MIMIC-IV for Correctness and Informativeness of non-low-resource codes, and the Acceptability of the discharge summaries. For the other metrics, a $\kappa>0$ was reached but never exceeded 0.4 (lower than moderate agreement). All mean scores are higher than 4. Hence, while the clinicians disagreed on the exact scores, they rated real discharge summaries positively. The disagreement may be due to clinicians being UK-based with significant differences in reporting style within the UK and the US (where MIMIC-IV is from).
For GPT-generated summaries, slight agreement was seen in acceptability, and fair agreement in the correctness of low-resource labels. All other metrics had poor agreement. Both correctness metrics scored above 4, with low-resource correctness surpassing 4.5—an encouraging outcome for our primary goal of generating low-resource code data. Mean informativeness in the low-resource scenario and authenticity of scores were at least 3. Once again, performance on the low-resource codes exceeded non-low-resource codes. Other metrics had $\mu$ scores above 2. Informativeness and authenticity for non-low-resource codes had a poor agreement ($\kappa<0$), while acceptability had some agreement with the lowest mean score of 2.225.
Table 4: Evaluator agreement ($\kappa$) and mean scores ($\mu$) for samples from MIMIC-IV (real), versus GPT-generated (synthetic) data.
Metrics $\kappa_{\text{real}}$ $\mu_{\text{real}}$ $\kappa_{\text{synthetic}}$ $\mu_{\text{synthetic}}$
Correctness – Non-Low-Resource -0.386 4.175 -0.163 4.375
Correctness – Low-Resource 0.043 4.350 0.206 4.525
Informativeness – Non-Low-Resource -0.155 4.550 -0.220 2.775
Informativeness – Low-Resource 0.241 4.675 -0.277 3.000
Authenticity – Patient 0.340 4.750 -0.078 3.150
Authenticity – Scenario 0.373 4.775 -0.333 2.250
Acceptability -0.056 4.550 0.035 2.225
While G generally produces correct notes, the clinical evaluators have identified several challenges in the generation of natural-looking clinical notes:
@@
G tends to copy all concepts mentioned in the prompt when generating a clinical note. While instruction-following is a desirable behaviour, excluding non-crucial details is essential when generating a natural-looking clinical note. Real clinical notes often omit irrelevant and less critical findings for brevity, particularly if the information is inferrable from surrounding contexts such as medications and treatments. For instance, G unnecessarily noted a normal BMI.
@@
G tends to use an overly technical and unnatural style when specifying diagnoses. For instance, G mentioned “anaemia, which was unspecified.” in the generated clinical note as it was prompted with “D64.9: Anemia, unspecified”. G also occasionally introduces vague phrases (e.g.“geriatric team provided supportive care, including behavioural interventions and medication management”) without further detail. This contrasts with the more streamlined language of real clinical notes.
3.3.3 GPT-3.5 lacks details when introducing supporting information
G tends to introduce crucial supporting information without sufficient details. For instance, G mentioned “Following a traumatic event” without further specification of the mentioned traumatic event, which is unacceptable in the clinical setting. This omission limits the overall informativeness of the patient’s medical context, potentially hindering the notes’ usability for a comprehensive view.
@@
G sometimes introduces improbable but possible details. For instance, G overemphasised the significance of a patient’s anxiety disorder regarding an episode of syncope and a subsequent facial fracture, which the clinicians consider unlikely.
@@
G does not effectively present diagnoses as interconnected, resulting in fragmented notes that lack coherence. The clinicians described G-generated clinical notes as collections of unrelated facts. For example, G presented complications of Type 1 diabetes mellitus separately without illustrating their relation. The lack of coherence between diagnoses may impede the plausibility of the clinical note and undermine the overall acceptability and usefulness of synthetic notes.
@@
G struggles to prioritise diagnoses based on clinical significance, which undermines the authenticity of the portrayed scenario. For example, G often places critical conditions on the same level as minor issues, such as impacted ear wax, cataracts, and conjunctival haemorrhage. Hence, we concluded that G struggles to effectively convey the relative clinical significance of certain diagnoses.
@@ 4 DISCUSSION, CONCLUSION AND FUTURE STUDIES
In this work, we have investigated the capability of GPT-3.5’s potential in augmenting ICD-10 coding for local neural models in low-resource scenarios. While overall performance dipped with synthetic data augmentation, filtered codeset evaluation showed improvements, especially in advanced models like LAAT and Multi-Res CNN. Error analysis indicated augmented models made fewer out-of-family predictions, with some shift to within-family errors (closer to the correct answer). Augmentation showed promise in improving the prediction of generated codes and their siblings. Zero-shot labels did not consistently benefit from the augmentation, emphasising the need for real data in augmentation success. However, a zero-shot code learned from the synthetic data was predicted correctly. The potential of LLM-generated discharge summaries should further be explored with different (e.g., local or specialised) LLMs, prompt engineering, and further supplementing the prompt with external knowledge (e.g., from ontologies).
In guided synthetic settings with ICD-10 descriptions, GPT-3.5 showed partial code identification ability displaying lesser over-/under-prediction tendencies than previously reported local models. It, however, struggled in the realistic scenario without in-prompt aid, performing below locally-trained models. Hence, the explored setup of producing a synthetic document based solely on the associated ICD codes is unsuitable for deployment in a clinical setting.
@@
One potential solution to generating synthetic discharge summaries involves restructuring the prompt to order diagnoses chronologically, providing their corresponding timestamps. This could guide LLMs in creating synthetic notes mirroring the chronological progression of a patient’s medical journey, enhancing coherence and prioritisation.
@@
5 Limitations
In this study, while the annotation experts are involved as co-authors, we ensured that they were independent from the development of the algorithms that involved the synthetic data. While the evaluation utilised few clinical experts (n=4), they provided sufficient expertise in evaluating the notes. The study was blinded with respect to the real/synthetic status of documents, but according to the experts the synthetic data differed from real enough to be distinguishable.
6 Data Availability
The synthetic discharge summary data generated as part of this study will be shared on reasonable request to the corresponding author upon presenting a certificate of completion of the CITI Data or Specimens Only Research course from the Collaborative Institutional Training Initiative program (https://physionet.org/about/citi-course/). The data has been accepted for publication and will be made available via PhysioNet [^10] https://doi.org/10.13026/bnc2-1a81.
@@ 9 Author Contributions
MF proposed exploring data augmentation for ICD coding with GPT-3.5 and developed the main ideas of the publication with APG, HD, AB, and BA. MF, APG, HD, AB, and BA contributed to the experimental design. MF designed and performed the document synthesis and the evaluation of GPT’s ICD-10 coding performance on GPT-generated data. APG evaluated GPT’s ICD-10 coding performance on MIMIC-IV data. MF, APG, HD, and LD designed the evaluation of discharge summaries by clinical staff. LD, SB, MH, RSP performed the clinical expert evaluation of the data, and MF and APG analysed its results. MF and APG wrote the manuscript. All authors participated in editing the manuscript. BA and AB supervised the project and provided feedback and input on the experiments and analyses.